OpenStudy (anonymous):

Solve the system by the elimination method. Check your work. x - 3y = 0 3y - 6 = 2x

OpenStudy (anonymous):

x - 3y = 0 -(1) 3y - 6 = 2x -(2) (1)+(2) (x - 3y) + (3y - 6) = 0+2x Can you solve it now?

OpenStudy (anonymous):

yup thanks!!